rishavraj (rishavraj):

here it is.... :)) @mathmath333

rishavraj (rishavraj):

if a+b+c = 3x Simplify \[(x - a)^3 + (x + b)^3 + (x-c)^3 \]

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & (x - a)^3 + (x + b)^3 + (x-c)^3=3[(x - a)(x + b)(x-c)]\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

next

rishavraj (rishavraj):

if \[x^2 + x + 1 = 0 \] evaluate \[x^3 + \frac{ 1 }{ x^3 } = 0\]

OpenStudy (mathmath333):

lol but is my answer correct

rishavraj (rishavraj):

@mathmath333 dont u think I would hav queestiond u if it was wrong -_-.... or u want me to write "Proud of u" lmoa....JK :P xD

OpenStudy (mathmath333):

ok but the way havent u asked the second question before

rishavraj (rishavraj):

really did I ...I dont think so

OpenStudy (mathmath333):

+2 is the answer

rishavraj (rishavraj):

ooops my I really did ask tht before -_- here if \[\frac{ a }{ b } + \frac{ b }{ a } = 1\] evaluate \[a^3 + b^3\] too easy though :P

rishavraj (rishavraj):

btw wo octagon wale ka soln post krna hai???

OpenStudy (mathmath333):

yes post the octagon that was interesting

rishavraj (rishavraj):

http://prntscr.com/am3uty see u get the fig ??

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & a^3 + b^3=0\hspace{.33em}\\~\\ \end{align}}\)

OpenStudy (mathmath333):

yes fig is easy

rishavraj (rishavraj):

so let the edge lengt of octagon be = a then \[a = 11 - 2x = 13 -2y = \sqrt{x^2 + y^2}\] and yeah thts correct

OpenStudy (mathmath333):

:) good one

rishavraj (rishavraj):

lol yeah...... do u know the trick to solve \[\sqrt{42+\sqrt{42+\sqrt{42+\sqrt{42........ \infty }}}}\]

OpenStudy (mathmath333):

7